In an arithmetic sequence 214 = 7 and a27 = 15. Determine a. ay (Type an integer or a simplified fraction.)
Answer #1

914=4, aay= 15 ana at(nald 914 = at (14-1)dan at 13 day Agya at (21-1) d=15 ata6d=15- solve eu 4 6 Q+ 13 d = y Q+a6d=15 -13d=